# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ Discussion
We formulated a cell-scale model for epithelial tissue dynamics based on the assumptions of adiabatic force balance in the regime of dominant cortical tensions. Experimental evidence [41, 27, 17] suggests that the cortical cytoskeleton which generates this tension behaves more like a muscle, where the tension level is set by active regulation, rather than a spring, where tension is a function of length. Subscribing to these assumptions has two important consequences: First, feedback loops are required to achieve and stabilize a force-balanced configuration [13]. This is in contrast to a network of springs where the relation between length and tension (the constitutive relationship) ensures that force balance is reached by minimization of the elastic energy. Second, force balance of cortical tensions does not fully constrain the tissue as it allows for isogonal (angle-preserving) deformations which change the lengths of interfaces without changing the angles at which they meet in vertices [13]. These isogonal degrees of freedom of the cortical network are governed by non-cortical mechanical stresses, arising e.g. from passive cell elasticity, representing cell-internal structures such as the nucleus [23, 24], microtubules, and intermediate filaments [25]. We find that this internal rigidity is essential to transduce cell intercalations into tissue-scale deformation. In the absence of cell resistance against deformation, intercalations are compensated by cell shape changes. However, if cell-internal elasticity becomes stronger than cortical tensions, it can resist changes in vertex angles and impede T1s and tissue flow (Fig. S11). Such a scenario may occur in certain genetic mutants, like the “kugelkern” (kuk) mutant of Drosophila, where the nucleus is stiffer [42]). Our findings suggest that epithelial tissue flows not like a fluid (where the shear modulus vanishes) but rather as a plastically deforming solid, whose remodeling is driven internally while resisting external forces. Epithelial tissue can thus be regarded as an active solid.
@@
Our model formulates a theory of active elasticity where the dominant (cortical) stresses are not governed by constitutive relationships. It builds explicitly on the geometric relation (duality) between tension space and real space afforded by the force balance condition. Stabilizing feedback mechanisms that maintain adiabatic force balance are implicit in our model as we constrain the tension dynamics to the space of force-balanced configurations (flat tension triangulations). Dynamics in tension space is formulated in terms of the tension triangulation geometry driven by local positive feedback. This feedback amplifies a weak initial tension anisotropy and thus drives cell shape dynamics that result in cell intercalations (T1 processes). Force balance provides the non-local coupling that allows for coordination of forces and cellular behaviors across the tissue. On the tissue scale, self-organized active T1s are oriented by global tension anisotropy and thus drive convergent-extension flow. As T1s drastically remodel tension geometry, they gradually degrade the orientational cue provided by initial tension anisotropy. Thus, the locally self-organized tissue flow is generically self-limiting. It arrests after a finite extent of convergent extension that depends on the initial degree of order in the cellular packing and the magnitude of initial tension anisotropy. This central finding suggests that cell geometry is a repository of morphogenetic information that may encode the final tissue shape.
In the companion paper [17], we showed that Drosophila germ-band extension is driven by internally generated forces (rather than by external pulling forces). The tissue-scale model presented here – describing local mechanical self-organization of cells – provides an elegant explanation of a range of experimental observations. Supplied with the experimentally observed initial packing order and tension anisotropy our model predicts the ca. two-fold elongation two-fold elongation of the germ band. In addition, our geometrically formulated tension dynamics model can be directly compared with experimental data on the cell scale [17]. To this end, we have introduced the LTC order parameter for the local configuration of tensions. Comparing the LTC time courses between experiments and simulations, we find excellent agreement, suggesting that our model reproduces the dynamics of germ band extension also on the cell scale. Specifically, our order parameter distinguishes two motifs of local tension configurations “tension cables” where multiple high tension interfaces meet in a vertex and “tension bridges”: high tension interfaces surrounded by low tension interfaces. Tension bridges are the elementary motif of the alternating pattern of tensions that choreographs active T1s across the tissue.
The dynamics in tension configuration space depends on the character of the positive tension feedback. Winner-takes-all feedback efficiently drives the local tension configuration toward the T1-threshold via the formation of tension bridges. By contrast, when feedback saturates at too low relative tension, it causes formation of tension cables, which have previously been suggested as a driver for convergent extension. However, our simulations and analysis of local tension configurations show that tension cables are inefficient at driving convergent extension, as adjacent interfaces “compete” to contract. Indeed, arrest of convergent extension due to the formation of tension cables is also observed in recent computational studies [10, 43]. Indeed, tension cables are often found to form at boundaries between distinct tissues and at segment boundaries where they prevent mixing between the adjacent tissues, i.e. prevent cell rearrangements [44, 45, 46, 47].
When tension cables contract, they lead to formation of “rosettes” where five or more cells meet in a single vertex. While it has been suggested that rosettes are important for epithelial convergent extension [41, 48], recent quantitative analysis of Drosophila gastrulation [17, 16, 20] shows that rosettes contribute significantly less to tissue extension than T1s and appear in conjunction with disorder which causes arrest of tissue flow. Our simulations corroborate these empirical findings. In a Drosophila mutant for abl, contraction of tension cables is impaired resulting in a reduction of rosette formation [49]. By contrast, T1s appear unaffected in these mutants, suggesting that isolated high-tension junctions contract normally. Our model offers a possible explanation of this puzzling finding: Deletion of abl might cause the positive tension feedback to saturate earlier thus leading to the formation of persistent tension cables. This hypothesis may guide future research into the molecular role of abl on the regulation of the junctional cytoskeleton.
Taken together, the findings discussed above show that tension bridges are key to drive efficient convergent-extension flow. In contrast to tension cables, which clearly stand out in microscopy images (e.g. of fluorescently labelled myosin), tension bridges are hard to spot as they rapidly contract. This might be a reason why the role of tension bridges has not been appreciated before. The tension configuration parameter introduced here facilitates a statistical analysis across many cells and allows one to distinguish different local tension dynamics (e.g. winner-takes-all vs saturating feedback).
@@
Our scenario for GBE is based on local self-organization driven by mechanical feedback. Self-organized T1s are facilitated by the initial order in the cell packing and are oriented by an initial global tension anisotropy but do not require cell-scale genetic instructions [33, 50]. We expect that the initial tension anisotropy, whose presence we confirmed experimentally in Ref. [17], is set up early on by anisotropic static “hoop” tension resulting from the internal turgor pressure in the embryo, and further reinforced by the dynamic effects of ventral furrow formation [14]. To create global tissue flow, local self-organization must be modulated by large-scale pre-patterning of cell behaviors. In the early Drosophila embryo, this is manifested in the dorso-ventral patterning system that specifies the tissues with different mechanical properties and modulates mechanical feedback loops [14].
In simulations on a cylindrical geometry without droso-ventral patterning, mimicking a Toll[RM9] mutant, active T1s occur in the absence of tissue flow. Instead, the cell rearrangements are compensated by cell shape deformations. Notably, these simulations also show that T1 resolution does not require external pulling forces or contractions of the cells’ apical area by “medial” myosin pulses [8]. Taken together, our model provides unified picture for Drosophila GBE that bridges the cell and tissue scales. The integration of bottom-up self-organization and top down genetic control emerges as a common theme in development [51].
@@
Our model predicts that disrupting the hexagonal packing of nuclei prior to cellularization will cause slower GBE. Interesting candidates to test this prediction are “nuclear fallout” mutants where some nuclei leave the blastoderm surface and thus introduce defects in the cellular packing [52]. Another option might be the transient and partial disruption of microtuble organization with small molecule inhibitors [53]. We expect that these experiments can be used to challenge and subsequently refine the model.
An important challenge for future work is to identify the (molecular) mechanisms that stabilize the force-balanced configuration on short timescales while driving controlled remodeling on long timescales. In our model, maintenance of force balance implicitly enforced by maintaining a flat tension triangulation (the geometric manifestation of global force balance). This subsumes the complex regulatory feedback loops operating in the cortical cytoskeleton. The underlying mechanics of the interplay of actin fibers, myosin motors, passive crosslinkers (such as spectrins [54]) and mechanical feedback mediators (such as $\alpha$-catenin [55]) remain poorly understood. We hope that the insights from the geometrical model and the quantification of local tension configurations [17] will help develop more fine-grained models that explicitly account for these details.
On the continuum level tissue flow can be described as a Stokes flow driven by an active stress field that captures the coarse grained distribution of myosin motors [1, 2, 56, 57]. Coupling of the active stress fields to the flow (by advection and flow alignment) generically gives rise to long-wavelength spontaneous flow instabilities [58, 59, 60] and active nematic turbulence [61]. Notably, we do not observe such an instability in our model. Instead, we find that tissue flow driven by local mechanical self-organization of cells is generically self-limiting due to the rapid loss of cell-scale order as cells rearrange. A similar “short-wavelength instability” is found in another recently model for self-organized T1s [10]. It is not captured by existing continuum models, suggesting that additional order parameters (such as the hexagon fraction and the LTC order parameter introduced here) will need to be taken into account. Interestingly, several biological systems exhibit persistent convergence–extension flows [62, 63] that appear to overcome the self-limiting nature of a locally self-organized mechanism. Such persistent flows may be accounted for by coupling tension dynamics to external fields [64] representing guiding cues such as planar cell polarity [65]. Overall, we anticipate that translating the intimate geometric relation between tension space triangulation and physical cell-array to the continuum level will be a fruitful direction for future research.
@@ Acknowledgements.
We thank Arthur Hernandez, Matthew Lefebvre, Noah Mitchell, Sebastian Streichan, and Eric Wieschaus for stimulating discussions and careful reading of the manuscript. We further thank Dinah Loerke and Jennifer Zallen for insightful feedback. FB acknowledges support of the GBMF post-doctoral fellowship (under grant #2919). NHC was supported by NIGMS R35-GM138203 and NSF PHY:1707973. BIS acknowledges support via NSF PHY:1707973 and NSF PHY:2210612.
@@ 1.2 Hexagonal packing and hexatic order
We quantify order of the cell arrangement in the tissue by two measures: (i) the fraction of six-sided cells, that one might loosely call “hexagonal packing order” and (ii) the hexatic order parameter measuring bond-orientational order. Hexagonal packing order is a topological quantity: it depends only on the neighborhood relations between cells but not on their exact shape. In contrast, hexatic order (defined below) is a geometric measure, that is sensitive to the cell shapes. Initially, the majority of cells have six neighbors (Fig. S1A, 0 min), and the density of topological defects, manifested as non-hexagonal cells, is relatively low. During VF invagination, the number of defects increases only slightly (25 min) since there are only very few intercalations [17]. During GBE, the number of defects increases significantly as a consequence of T1s. Towards the end of GBE, the distribution of cell-coordination number approaches that of a random Voronoi tessellation seeded with a Ginibre random point process (see Fig. S1B and B’).
Figure S1: Coordination number and hexatic order of cells. A Initially (0 min), the majority of cells are hexagonal (light gray). During GBE (25–50 min), topological defects proliferate, resulting in increasing numbers of non-hexagonal cells. Semitransparent gray overlay marks cells that invaginate. B and B’ Fraction of hexagonal cells as a function of time (B) and histogram of cell-coordination number (B’) in the lateral ectoderm. The histograms show pooled data from the time periods 0–15 min and 45–50 min highlighted in B. Lines indicate the coordination number distributions for Voronoi tessellations generated from three different random point processes: hard disk (packing fraction 0.54 near the theoretical maximum for sequential placement [66]), Ginibre [67], and Poisson. C Hexatic order is low, and shows no long-range correlation. It further decreases as GBE progresses. Hue and brightness code for the phase ($\mathrm{arg}\,\psi_{6}$) and magnitude ($|\psi_{6}|$) of the hexatic order parameter, respectively.
@@
The hexatic order parameter (also called “bond orientational order parameter” [68]) for a cell with $n$ vertices is defined as
@@
where $\theta_{k}$ is the angle between the DV axis and the vector pointing from the cell’s centroid to vertex $k$. For a cell with regular hexagonal shape the magnitude of $\psi_{6}$ reaches its maximum $|\psi_{6}|=1$. The phase $\mathrm{arg}\,\psi_{6}$ indicates the orientation of the hexagon. In contrast to the coordination number, which is a purely topological measure, the hexatic order parameter depends on the cell shape. In line with previous findings from confocal microscopy of the Drosophila germ band [69], we find that the hexatic order parameter to be low in magnitude and exhibit no long-range correlations as is apparent from Fig. S1C. Effectively, the presence of non-hexagonal cells act as obstructions to long range correlations of geometric order. We quantify the range of correlations in hexatic order by coarse graining over patches of cells with different radii (measured in the by the neighborhood degree). Before the onset of GBE, the coarse grained order parameter decays as a power law of the patch radius law with an exponent that close to $-0.75$, a value found also for MDCK cells and in simulations using multiphase-field models [70]. (We find the same result using the distance-weighted hexatic order parameter introduced in Ref. [70].) At the end of GBE, the decay exponent is smaller than $-1$, indicating a complete lack of correlations in hexatic order between neighboring cells. This is expected from the high number of non-hexagonal cells at this stage (Fig. S1B, 50 min).
@@ 1.3 Local and regional tension inference
Figure S2: Comparison of local vs regional tension inference. A Local tension inference at a single vertex through complementarity of angles at a single vertex (left) and in the corresponding tension triangle (right). The relative tensions (edge lengths of the tension triangle) are found via the law of sines. B Regional tension inference for a patch of three cells. The interfaces (and corresponding tensions) shared between them are highlighted by bold lines. The triangulation property of the tensions yields one additional constraint for each cell because for each cell there is a loop in the tension triangulation which has to close. One such loop is highlighted. These additional constraints make the regional inference more robust. C Comparison of the LTC distributions for regional and local tension inference in the Drosophila germ band. Distributions are from data collected over the indicated time intervals. Germ band elongation starts at ca. 25 min. The last row shows single-vertex inference based on vertex positions smoothed using a moving average (1.25 min interval). D Mean anisotropy magnitude (top) and LTC phase (bottom) for the three inference variants. Note the systematically lower LTC phase, i.e. lower fraction of tension bridges, found using single-vertex inference.
To quantify the local tension configurations in experimental data from the Drosophila germ band, we performed tension inference based on segmented cell outlines. In the companion paper [17], tension inference was done locally, directly relating the angles at each vertex to the relative cortical tension via the law of sines (see Fig. S2A). While this method is conceptually and computationally simple, it is sensitive to observational and dynamical noise in the angles. A more robust approach is to perform tension inference in an extended region, which makes the inference problem overconstrained [13]. In particular, there is one additional constraint per cell because the tension triangles for each cell have to fit together in force balance as illustrated in Fig. S2B. The overconstrained inference returns tensions for the closest force-balanced configuration compatible with the observed angles and thus removes small deviations from tensional force balance due to pressure differentials, short-time fluctuations, and observational noise. (An alternative method to remove observational noise is to apply a moving average on the vertex positions before inferring tensions.) Since in our model cortical tensions are always in exact tensional force balance, we use this overconstrained inference to compare to simulations. Specifically, we use tension inference on all interfaces of the three cells which meet at a given vertex to determine the local tension configuration parameter at that vertex. Figure S2C shows the LTC distributions obtained from local and regional inference. The LTC distributions are qualitatively very similar but they differ quantitatively. In particular, the LTC phase $\tilde{\psi}$ has the same qualitative trend with a transient increase before the onset of T1s but is systematically lower for regional inference.
Figure S3: LTC distribution for the hard-disk Delaunay family of random triangulation. A  Median LTC parameter magnitude and phase vs hard disk packing fraction. B, C LTC histograms for two different packing fractions shows significant differences in the LTC distribution for different $\rho$. Note that topological statistics such as coordination numbers are much less sensitive to $\rho$. LTC distributions computed from cell arrays with $4000$ cells.
Figure S4: Illustration of LTC parameters as a function of internal triangle angles. The three angles $\alpha,\beta,\gamma$ of a triangle lie in a simplex defined by $\alpha,\beta,\gamma\geq 0$ and $\alpha+\beta+\gamma=\pi$. The simplex has a 6-fold symmetry due to permutations of the angles; its fundamental domain is mapped into the LTC space formed by $\tilde{s},\tilde{\psi}$ defined in Eq. (8). The Delaunay threshold $\max\{\alpha,\beta,\gamma\}\leq\pi$ is indicated in gray.

### v3

@@ Discussion
Our model provides a theory of active elasticity based on the geometric relation (duality) between tension space and real space afforded by the force balance condition. Stabilizing feedback mechanisms that maintain adiabatic force balance are implicit in our model as we constrain the tension dynamics to the space of force-balanced configurations (flat tension triangulations). Dynamics in tension space is driven by local positive feedback. This feedback amplifies a weak initial tension anisotropy and thus drives cell shape dynamics that result in cell rearrangements (T1 processes). Notably, while active T1s are initiated by positive tension feedback, we find that their resolution is through passive relaxation of tension and does not require additional active ingredients as previously suggested [Collinet.etal2015]. Force balance provides the non-local coupling that allows for the coordination of forces and cellular behaviors across the tissue. On the tissue scale, self-organized active T1s are oriented by global tension anisotropy and thus act coherently to drive convergent-extension flow. As T1s drastically remodel tension geometry, they gradually degrade the orientational cue provided by initial tension anisotropy. (Note that this degradation of order happens locally through T1 transitions; its relation to long-wavelength instabilities of active nematics in which there are no explicit topological transitions is an interesting avenue for future research.) Thus, the locally self-organized tissue flow is generically self-limiting. It arrests after a finite extent of convergent extension (CE) that depends on the initial degree of order in the cellular packing and the magnitude of initial tension anisotropy. This central finding suggests that cell geometry is a repository of morphogenetic information that may encode the final tissue shape.
Importantly, we show that T1s are controlled by local configuration of tensions as quantified by the LTC order parameter that links the locally hexatic space of the tension triangulation (representing force balance) with the nematic nature of deviatoric stress and strain in physical space. In contrast to previous work on “hexanematic” order in tissues [Armengol-Collado.etal2023, Armengol-Collado.etal2024], LTC order is defined in tension space, not in physical space, and does not require (quasi-) long-range hexatic order. Moreover, the emergence of this order is not driven by a free energy but by active feedback acting on tensions as discussed further below.
Mechanically self-organized tissue dynamics provides an elegant explanation for the arrest of Drosophila GBE after about two-fold elongation [Irvine.Wieschaus1994], while predicting its dependence on the initial tension anisotropy [Streichan.etal2018, Gustafson.etal2022]. In the companion paper [Brauns.etal2024elife], we used tension inference and LTC analysis to re-examine live imaging data on Drosophila germ-band extension (GBE) [Stern.etal2022] showing that it is driven predominantly by the internally generated forces in the lateral ectoderm of the embryo. Quantitative analysis of imaging data confirms that the Drosophila ectoderm starts ordered and becomes disordered as T1s proceed during CE.
@@
Notably, while our model does not require cell-scale genetic instructions to generate local tension anisotropy[Irvine.Wieschaus1994, Pare.etal2014], genetic patterning on the scale of the embryo is essential for the coordination and stability of the global flow. In the early Drosophila embryo, this is manifested in the dorso-ventral patterning system that specifies the tissues with different mechanical properties and modulates mechanical feedback loops [Gustafson.etal2022]. We expect that the initial tension anisotropy, reported in [Brauns.etal2024elife], is set up by anisotropic static “hoop” tension resulting from turgor pressure inside the embryo, and further reinforced by the dynamic effects of ventral furrow formation [Gustafson.etal2022]. An important challenge for future work is to identify the (molecular) mechanisms of both positive and negative feedback circuits, the latter stabilizing the force-balanced configuration on short timescales while the former, driving controlled remodeling on long timescales.
@@
Our model predicts that disrupting the hexagonal packing of nuclei prior to cellularization will cause slower GBE. Interesting candidates to test this prediction are “nuclear fallout” mutants, where some nuclei leave the blastoderm surface and thus introduce defects in the cellular packing   citeRothwell.etal1998. Another option might be the transient and partial disruption of microtubule organization with small molecule inhibitors [Kanesaki.etal2011]. We expect that these experiments can be used to challenge and subsequently refine the model.
Comparing the LTC time courses between experiments and simulations, we find an excellent agreement, suggesting that positive feedback-driven local tension dynamics can explain cell-scale behavior during GBE. The dynamics in tension configuration space depends on the character of the positive tension feedback. Winner-takes-all feedback efficiently drives the local tension configuration toward the T1 threshold via the formation of tension bridges. By contrast, when feedback saturates at too low relative tension, it causes the formation of tension cables, which had previously been suggested as a driver for CE. However, our simulations and analysis of local tension configurations show that tension cables are inefficient at driving CE, as adjacent interfaces “compete” to contract. Indeed, the arrest of CE due to the formation of tension cables is also observed in recent computational studies [Sknepnek.etal2023, Rozman.etal2023]. When tension cables contract, they lead to the formation of “rosettes” where five or more cells meet in a single vertex [Zallen.Zallen2004, Blankenship.etal2006]. In a Drosophila mutant for abl, contraction of tension cables is impaired resulting in a reduction of rosette formation [Tamada.etal2012]. By contrast, T1s appear unaffected in these mutants, suggesting that isolated high-tension junctions contract normally. Our model offers a possible explanation of this puzzling finding: Deletion of abl might cause the positive tension feedback to saturate earlier thus leading to the formation of persistent tension cables. Taken together, the findings discussed above show that tension bridges are key to driving efficient convergent-extension flow. The LTC order parameter introduced here facilitates statistical analysis across many cells and allows one to distinguish different regimes of local tension dynamics.
Because force balance geometry does not uniquely define the shape of cells – on account of isogonal degrees of freedom – the latter play an important role in defining tissue dynamics. Isogonal soft modes account for tissue deformation under constant cortical tensions and are controlled by non-cortical mechanical stresses, arising e.g. from passive cell elasticity due to cell-internal structures (e.g. nucleus [Grosser.etal2021, Kim.etal2022nuclear], microtubules, and intermediate filaments [Pensalfini.etal2023]). We find that internal rigidity is essential to transduce cell intercalations into tissue-scale deformation against resistance from adjacent tissues. In the absence of cell resistance against deformation, intercalations are compensated by cell shape changes. However, if cell-internal elasticity becomes stronger than cortical tensions, it can resist changes in vertex angles and impede T1s and tissue flow (Fig. S11). Such a scenario may occur in certain genetic mutants, like the “kugelkern” (kuk) mutant of Drosophila, where the nucleus is stiffer [Brandt.etal2006]).
Overall, our findings suggest that epithelial tissue flows not like a fluid (where the shear modulus vanishes) but rather as a plastically deforming solid, whose remodeling is driven internally, while resisting external forces. Epithelial tissue can thus be regarded as an active solid. More generally, tissue mechanics dominated by cortical tensions that are controlled by feedback mechanisms call for a formulation of continuum mechanics that does not rely on a constitutive stress-strain relationship.
@@ Acknowledgements.
We thank Arthur Hernandez, Matthew Lefebvre, Noah Mitchell, Sebastian Streichan, and Eric Wieschaus for stimulating discussions. We further thank Dinah Loerke and Jennifer Zallen for their insightful feedback. FB acknowledges support of the GBMF post-doctoral fellowship (under grant #2919). NHC was supported by NIGMS R35-GM138203 and NSF PHY:1707973. BIS acknowledges support via NSF PHY:1707973 and NSF PHY:2210612.
@@ 1.2 Hexagonal packing and hexatic order
We quantify the order of the cell packing in the tissue by two measures: (i) the fraction of six-sided cells which one might loosely call “hexagonal packing order” and (ii) the hexatic order parameter measuring bond-orientational order. Hexagonal packing order is a topological quantity: it depends only on the neighborhood relations between cells but not on their exact shape. In contrast, hexatic order (defined below) is a geometric measure sensitive to the cell shapes. Initially, the majority of cells have six neighbors (Fig. S1A, 0 min), and the density of topological defects, manifested as non-hexagonal cells, is relatively low. During VF invagination, the number of defects increases only slightly (25 min) since there are only very few intercalations [Brauns.etal2024elife]. During GBE, the number of defects increases significantly due to T1s. Towards the end of GBE, the distribution of cell-coordination number approaches that of a random Voronoi tessellation seeded with a Ginibre random point process (see Fig. S1B and B’).
@@
The hexatic order parameter (also called “bond orientational order parameter” [Halperin.Nelson1978]) for a cell with $n$ vertices is defined as
@@
where $\theta_{k}$ is the angle between the DV axis and the vector pointing from the cell’s centroid to vertex $k$. For a cell with a regular hexagonal shape, the magnitude of $\psi_{6}$ reaches its maximum $|\psi_{6}|=1$. The phase $\mathrm{arg}\,\psi_{6}$ indicates the orientation of the hexagon. In contrast to the coordination number, which is a purely topological measure, the hexatic order parameter depends on the cell shape. In line with previous findings from confocal microscopy of the Drosophila germ band [Zallen.Zallen2004], we find that the hexatic order parameter is low in magnitude and exhibits no long-range correlations as is apparent from Fig. S1C. Effectively, the presence of non-hexagonal cells acts as an obstruction to long-range correlations of geometric order. We quantify the range of correlations in hexatic order by coarse-graining over patches of cells with different radii (measured by the neighborhood degree). Before the onset of GBE, the coarse-grained order parameter decays as a power law of the patch radius law with an exponent that is close to $-0.75$, a value found also for MDCK cells and in simulations using multiphase-field models [Armengol-Collado.etal2023]. (We find the same result using the distance-weighted hexatic order parameter introduced in Ref. [Armengol-Collado.etal2023].) At the end of GBE, the decay exponent is smaller than $-1$, indicating a complete lack of correlations in hexatic order between neighboring cells. This is expected from the high number of non-hexagonal cells at this stage (Fig. S1B, 50 min).
@@ 1.3 Local and regional tension inference
To quantify the local tension configurations in experimental data from the Drosophila germ band, we performed tension inference based on segmented cell outlines. In the companion paper [Brauns.etal2024elife], tension inference was done locally, directly relating the angles at each vertex to the relative cortical tension via the law of sines (see Fig. S2A). While this method is conceptually and computationally simple, it is sensitive to observational and dynamical noise in the angles. A more robust approach is to perform tension inference in an extended region, which makes the inference problem overconstrained [Noll.etal2017]. In particular, there is one additional constraint per cell because the tension triangles for each cell have to fit together in force balance as illustrated in Fig. S2B. The overconstrained inference returns tensions for the closest force-balanced configuration compatible with the observed angles and thus removes small deviations from tensional force balance due to pressure differentials, short-time fluctuations, and observational noise. (An alternative method to remove observational noise is to apply a moving average on the vertex positions before inferring tensions.) Since in our model cortical tensions are always in exact tensional force balance, we use this overconstrained inference to compare to simulations. Specifically, we use tension inference on all interfaces of the three cells that meet at a given vertex to determine the local tension configuration parameter at that vertex. Figure S2C shows the LTC distributions obtained from local and regional inference. The LTC distributions are qualitatively very similar but they differ quantitatively. In particular, the LTC phase $\tilde{\psi}$ has the same qualitative trend with a transient increase before the onset of T1s but is systematically lower for regional inference.